OpenStudy (trisarahtops):

A snowball is melting at the rate of 2 inches per hour. How fast is the surface area of the snowball changing at the instant the snowball has a radius of 3 inches? −16π in2/hour −24π in2/hour −32π in2/hour −48π in2/hour (I think it's this one)

OpenStudy (trisarahtops):

@paki

OpenStudy (er.mohd.amir):

surface area=4*pi*r*r given dr/dt=2 inch/h

OpenStudy (er.mohd.amir):

so ds/dt=4*pi*(2*r)*(dr/dt)

OpenStudy (trisarahtops):

okay so 8*pi*(3)(-2) right?

OpenStudy (trisarahtops):

or 4*pi*(2*3)*2

OpenStudy (trisarahtops):

hello?

OpenStudy (er.mohd.amir):

yes ds/dt=-2 ans -48*pi

OpenStudy (trisarahtops):

Thank you ;)